20 - 1/5d = 3/10d + 16
Answer by checkley77(12844)

20 - 1/5d = 3/10d + 16
20-d/5=3d/10+16
-d/5-3d/10=16-20
(-2d-3d)/10=-4 cross multiply.
-5d=-40
d=-40/-5
d=8 ans.
Proof:
20-1/5*8=3/10*8+16
20-8/5=24/10+16
-8/5-24/10=16-20
(-8*2-24)/10=-4
(-16-24)/10=-4
-40/10=-4
-4=-4